# Policy Space Response Oracles: A Survey

### v1

@@ Abstract
In game theory, a game refers to a model of interaction among rational decision-makers or players, making choices with the goal of achieving their individual objectives. Understanding their behavior in games is often referred to as game reasoning. This survey provides a comprehensive overview of a fast-developing game-reasoning framework for large games, known as Policy Space Response Oracles (PSRO). We first motivate PSRO, provide historical context, and position PSRO within game-reasoning approaches. We then focus on the strategy exploration problem for PSRO, the challenge of assembling an effective strategy portfolio for modeling the underlying game with minimum computational cost. We also survey current research directions for enhancing the efficiency of PSRO, and explore the applications of PSRO across various domains. We conclude by discussing open questions and future research.
@@ 1 Introduction
In recent decades, the exploration of multiagent systems has been a central focus in Artificial Intelligence (AI) research. A multiagent system, often referred to as a game, comprises multiple decision-making agents that interact within a shared environment. To understand the strategic behavior among these agents – where the optimal behavior of one agent depends on the behavior of others – game theory provides a mathematical framework that defines behavioral stability through solution concepts like the Nash equilibrium (NE). To identify such solutions, various equilibrium computation approaches have been developed von Stengel (2002); Savani and Turocy (2023), either to enumerate all equilibria (see the work by Avis et al. Avis et al. (2010) for bimatrix games), or to find a single sample equilibrium (e.g., with the Lemke-Howson algorithm for a bimatrix game, or Linear Programming for a zero-sum matrix game von Stengel (2002)).
@@
As the size of the game (i.e., the number of players and strategies) grows, the computational feasibility of enumeration diminishes, and one tends to focus on finding a sample equilibrium. Unlike for general-sum games, for zero-sum games, a sample equilibrium already provides valuable insights into effective strategic play, for example, by revealing the game’s (unique) value, which can be achieved with a strong strategy irrespective of the strategy of the other player. However, even in zero-sum settings, many games that arise from practical applications are too large to solve for a sample equilibrium even with polynomial-time methods for Linear Programming – these games can be extremely large that they are infeasible to be explicitly represented as a payoff matrix in practice. It is these huge games that are our primary focus in this survey.
@@
As an alternative to traditional equilibrium computation methods, to reason about such huge games, a wide range of learning methods have been applied (with one of the most prominent approaches being multiagent Reinforcement Learning (RL) Albrecht et al. (2024)). While learning approaches have significantly contributed to the development of intelligent agents, they face many inherent challenges in games. For example, independent learning across agents can render the environment non-stationary, which is a challenge for convergence as each individual learner faces a potentially moving target. Another challenge is non-transitivity of a game, where there is not a clear notion of “better” strategy for an agent, and thus effective learning requires the learning scheme to maintain a population of strategies for each agent.
@@
Against this backdrop, the Policy Space Response Oracles (PSRO) framework Lanctot et al. (2017) emerged as a natural blend that combines traditional game-theoretic equilibrium computation with learning. In PSRO, a key concept is an empirical game model, which acts as an approximation of the underlying full game. Game models are induced from simulations run over combinations of a particular set of strategies. This set of strategies is typically much smaller than the full game, and thus game models are feasible to analyze with traditional equilibrium computation methods. PSRO alternates between the analysis of the current game model, defining a new learning target, and game model refinement by including the new strategies generated via learning.
@@
As a general solver for large-scale games, PSRO has been successfully applied to a wide range of game types and diverse application domains, from mechanism design Zhang et al. (2023) to robust RL Liang et al. (2023). Numerous PSRO variants have been developed, each tailored to leverage the specific characteristics of the underlying games. As a notable example of its success, algorithms inspired by PSRO have reached state-of-the-art performance in large-scale games such as Barrage Stratego McAleer et al. (2020), and in StarCraft Vinyals et al. (2019), where they have convincingly outperformed human experts and prior AI systems.
@@
While PSRO and its variants have been covered to some extent in existing multiagent learning surveys (e.g., Yang and Wang (2020); Long et al. (2023)), a dedicated survey on PSRO like this one has been lacking. In this survey, we reflect on the historical development of PSRO, which arose from different research communities, and we place PSRO within the space of game reasoning approaches. We then present the latest developments on PSRO, highlighting both current and future research directions.
@@ 2 The PSRO Framework
The PSRO framework incorporates a synthesis of ideas originating from distinct research communities. Within the planning community, McMahan et al. McMahan et al. (2003) laid the groundwork by formulating robust planning in Markov Decision Processes as a zero-sum game characterized by a vast strategy space. Drawing inspiration from Bender’s decomposition in optimization Benders (1962), they introduced the Double Oracle (DO) algorithm. DO is an iterative algorithm for solving games with a finite number of strategies. DO maintains a restricted version of the full game and iteratively expands the restricted game by adding best responses to the current equilibrium. When DO terminates, no player can deviate unilaterally to gain extra payoff and therefore the equilibrium in the current restricted game is an NE of the full game. In finite games, DO is guaranteed to converge to NE, though the restricted game could include all strategies of the full game in the worst case.
@@
Concurrently, similar methods were being developed from a different perspective in the co-evolution research community. Co-evolutionary methods evolve multiple populations of (different) species in parallel. In this setting, there is no explicit fitness function, but the fitness of a population member depends on how well it interacts with members of other populations. Traditional challenges for co-evolution include the presence of intransitive cycles in traits, and the related problem of forgetting a trait that seems not useful at one point in the evolutionary process but later becomes useful again. The community explored memory mechanisms and game formulations where the co-evolutionary process was setup to discover NE as mixtures of traits Angeline et al. (1993); Popovici et al. (2012). To overcome the mentioned obstacles of intransitive cycles and forgetting traits, various archival structures were devised to facilitate monotonic convergence towards a range of game-theoretic solution concepts. For example, the “Nash memory” for symmetric games Ficici and Pollack (2003) identifies equilibrium strategies within a discovered restricted game. This was extended to asymmetric games via the Parallel Nash Memory Oliehoek et al. (2006), broadening the standard DO framework to accommodate alternative oracle types beyond the best response.
@@
The PSRO framework Lanctot et al. (2017) was introduced in a paper titled “A Unified Game-Theoretic Approach to Multiagent Reinforcement Learning”, bringing many of the mentioned ideas from the planning, multiagent RL, and co-evolution communities together in a unified framework, which also covers classical game-theoretic learning dynamics such as Fictitious Play Brown (1951). Technically, PSRO generalizes DO by introducing the concept of meta-strategy solver (MSS), which extracts a profile from the current restricted game as the next best-response target (see Figure 1). This enables the best-response target to extend beyond NE, transforming PSRO into a versatile framework capable of generalizing various classical and modern game-theoretic algorithms. For simplicity, we often refer to a solution concept as the MSS that computes it when meaning is clear from the context. For example, we may say “Nash” to mean the MSS that computes NE. Moreover, the original version of PSRO specified deep RL as the best-response oracle [^1] PSRO actually works with any form of (approximate) response oracles, including search, planning, and evolutionary algorithms etc. For example, Li and Wellman Li and Wellman (2021) applied natural evolution strategies Wierstra et al. (2014) as the best response oracle for PSRO., enabling best-response computation in environments with a large number of states and actions. Finally, the payoffs of profiles in PSRO are estimated through simulation, and therefore a restricted game with estimated payoffs is usually called an empirical game. This idea of combining game-theoretic analysis with simulation in PSRO has been known as Empirical Game-Theoretic Analysis (EGTA) Wellman (2006), which provides a broad toolbox of techniques for game reasoning with game models based on simulation data. In an early EGTA work, Schvartzman and Wellman Schvartzman and Wellman (2009) deployed tabular RL as a best-response oracle (at a time when deep RL did not exist yet) and NE of the empirical game as a best-response target for strategy generation. PSRO extends tabular RL to deep RL and can be viewed as an instance of EGTA with strategies generated by iteratively best responding to MSS solutions using deep RL.
@@ 2.1 The Framework
A normal-form (aka strategic-form) representation of the full game $\mathcal{G}=(N,(S_{i}),(u_{i}))$ is a tuple, where $N$ is a finite set of players, each with a non-empty set of strategies $S_{i}$ and a utility function $u_{i}:\Pi_{j\in N}S_{j}\rightarrow\mathbb{R}$. A restricted game $\mathcal{G}_{S\downarrow X}$ is a projection of the full game $\mathcal{G}$, with players choosing from restricted strategy sets $X_{i}\subseteq S_{i}$. An empirical game $\hat{\mathcal{G}}_{S\downarrow X}=(N,(X_{i}),(\hat{u}_{i}))$ is a restricted game with simulated payoffs, where $\hat{u}_{i}$ represents the utility function estimated over the restricted strategy space $X$.
@@ 2.2 Strategy Exploration in PSRO
In essence, game-theoretic analysis in PSRO is performed by reasoning about empirical game models, induced from simulations run over combinations of strategies. To construct a feasible and effective game model for game analysis, the selection of the strategies is pivotal. In particular, a game model is expected to contain a much smaller number of strategies than the full game for representation tractability yet still maintain the key strategic information of the full game Balduzzi et al. (2018). This challenge of game model construction is described as the strategy exploration problem Jordan et al. (2010), which is the main research focus for developing PSRO methods. The goal of strategy exploration is to assemble an effective strategy portfolio for a game model with minimum computational cost (i.e., with the fewest strategies required). In PSRO, strategy exploration can be controlled by setting MSSs and ROs, which have a coupled impact; we refer to the joint choice as an MSS-RO combination.
@@ 2.3 Organization of the Survey
As discussed in Section 2.2, a key design question for PSRO revolves around how to conduct effective strategy exploration given a specific goal. Although the performance of strategy exploration depends on the interplay between the chosen MSSs and ROs, existing literature predominantly focuses on setting either MSSs or ROs independently. Therefore, we organize our discussion on research directions and corresponding PSRO variants by first discussing setting MSSs and ROs independently (Sections 3 and 4 respectively). We then, in Section 5, discuss works that have investigated the joint choice of MSS-RO combination, before moving on to discuss in Section 6 how best to evaluate the effectiveness of such choices for strategy exploration. We then discuss research on improving the efficiency of PSRO (Section 7), explore applications of PSRO (Section 8), and conclude with open questions and future research directions (Section 9).
@@ 3 Strategy Exploration via MSS
In prior works, setting MSSs was a primary way to control strategy exploration. In this section, we discuss these works, their motivations, and their efficacy for strategy exploration.
3.1 PSRO with Normal-Form Empirical Games
In the standard PSRO framework, an empirical game is represented in normal form. While the simulations typically unfold through sequential observations and decisions over time, the empirical game abstracts away this temporal structure. This section deals with research questions and PSRO variants based on the normal-form representation.
@@ 3.1 PSRO with Normal-Form Empirical Games
In PSRO, the most common MSS target is NE, which can be computed by various game-theoretic methods based on the normal-form empirical game. PSRO with NE is essentially DO with deep RL for computing (approximate) best responses. Therefore, PSRO with NE inherits the convergence property of DO, given mild assumptions about the quality of the best responses: In finite games, as long as beneficial deviations can always be found with non-zero probability, PSRO with NE as MSS target will converge to an NE in the limit.
@@
Despite its convergence guarantee in the limit, achieving exact convergence in large games is often unattainable due to constraints such as limited computational resources. Consequently, many prior works studying strategy exploration in large games revolves around the development of new algorithms that exhibit strong empirical performance (e.g., rapid convergence in terms of regret within a small number of PSRO iterations). The overfitting problem has two meanings for strategy exploration. First, it means that strategy exploration may overfit to the NE of the empirical game, which may not be an effective best-response target due to the limited information in the empirical game, potentially yielding ineffective strategy exploration. Second, it means overfitting to a specific equilibrium in general games (e.g., games with more than two players or general-sum games) without sufficiently exploring the whole strategy space. Note that the overfitting can be a problem for any solution concept; we discuss it in the context of NE specifically due NE’s prominence as the MSS target in the literature.
@@
To address the overfitting problem, Lanctot et al. Lanctot et al. (2017) proposed an MSS, called Projected Replicator Dynamics (PRD), an adaptation of traditional replicator dynamics Taylor and Jonker (1978). PRD ensures a probability lower bound for selecting each strategy in the empirical game, allowing the new best response to train against not only strategies in the equilibrium support but also those outside the support. PRD can be viewed as a form of regularization to prevent overfitting the best response to the exact NE of the empirical game. Due to the diverse training targets, PRD also improves the stability of the new strategy.
Building on the concept of regularization, subsequent research has focused on designing MSSs that prevent overfitting by effectively regularizing the target profile. For instance, Wang et al. Wang et al. (2019) proposed an MSS that combines NE with a uniform distribution as the target profile, enabling the best response to an NE strategy mixed with exploration elements. Wright et al. Wright et al. (2019) developed a history-aware approach with a best-response (BR) target mixed with previous targets. Online double oracle Dinh et al. (2022) integrated PSRO with online learning to selected the BR target.
Wang and Wellman Wang and Wellman (2023b) addressed the overfitting problem by adopting an explicit view of regularization and introduced Regularized Replicator Dynamics (RRD), an MSS variant that truncates the NE search process in intermediate game models based on a regret criterion. Specifically, RRD computes the BR target by running RD in the empirical game, stopping once the regret of the current profile with respect to the empirical game meets a specified regret threshold. The regret criterion enables RRD to support direct control of the degree of regularization and can be adjusted to fit a specific game.

### v2

@@ Abstract
Game theory provides a mathematical way to study the interaction between multiple decision makers. However, classical game-theoretic analysis is limited in scalability due to the large number of strategies, precluding direct application to more complex scenarios. This survey provides a comprehensive overview of a framework for large games, known as Policy Space Response Oracles (PSRO), which holds promise to improve scalability by focusing attention on sufficient subsets of strategies. We first motivate PSRO and provide historical context. We then focus on the strategy exploration problem for PSRO: the challenge of assembling effective subsets of strategies that still represent the original game well with minimum computational cost. We survey current research directions for enhancing the efficiency of PSRO, and explore the applications of PSRO across various domains. We conclude by discussing open questions and future research.
@@ 1 Introduction
In recent decades, the exploration of multiagent systems has been a central focus in Artificial Intelligence (AI) research. A multiagent system, often referred to as a game, comprises multiple decision-making agents that interact within a shared environment. To understand strategic behavior among these agents – where the optimal behavior of one agent depends on the behavior of others – game theory provides a mathematical framework that defines behavioral stability through solution concepts like the Nash equilibrium (NE). To identify such solutions, various equilibrium computation approaches have been developed von Stengel (2002); Savani and Turocy (2024), either to enumerate all equilibria (see the work by Avis et al. Avis et al. (2010) for bimatrix games), or to find a single sample equilibrium (e.g., with the Lemke-Howson algorithm for a bimatrix game, or Linear Programming for a zero-sum matrix game von Stengel (2002)).
@@
As the size of the game (i.e., the number of players and strategies) grows, the computational feasibility of enumeration diminishes, and one tends to focus on finding a sample equilibrium. In the special case of two-player zero-sum games (i.e., settings where two players strictly compete), a sample equilibrium already provides valuable insights: it reveals the game’s unique value, which is the payoff that the first player can guarantee to obtain by playing a sufficiently strong (equilibrium) strategy irrespective of the strategy of the other player. However, even in zero-sum settings, many games that arise from practical applications are too large, and computing a sample equilibrium (even with polynomial-time methods for the resulting linear program) is infeasible. It is these huge games that are our primary focus in this survey.
@@
As an alternative to traditional equilibrium computation methods, to reason about such huge games, a wide range of learning methods have been applied. Applying learning methods to games is known as multiagent learning Shoham et al. (2007), with one of the most prominent approaches being multiagent Reinforcement Learning (RL) Albrecht et al. (2024). Compared to traditional methods, learning methods reduce the need to represent the entire game and create intelligent agents by exploring the game interactively. While learning approaches have significantly contributed to the development of intelligent agents, they face many inherent challenges in games. For example, independent learning across agents can render the environment non-stationary, which is a challenge for convergence as each individual learner faces a potentially moving target Tuyls and Weiss (2012). Another challenge is non-transitivity of a game, where there is not a clear notion of “better” strategy for an agent, and thus effective learning requires the learning scheme to maintain a population of strategies for each agent. Such non-transitivity exists ubiquitously in various games Czarnecki et al. (2020); Sanjaya et al. (2022); Li et al. (2023b).
@@
Against this backdrop, the Policy Space Response Oracles (PSRO) framework Lanctot et al. (2017) emerged as a natural combination of traditional game-theoretic equilibrium computation with learning. In PSRO, a key concept is a restricted game with estimated payoffs [^1] To be precise, normally a restricted game constrains the strategy space (relative to the full game), but the payoffs are “exact”, rather than being estimates. By contrast, in practice, PSRO implementations would typically estimate these payoffs through simulation (with stochasticity coming potentially from the environment). A restricted game with estimated payoffs would normally be termed an empirical game Wellman (2006) or meta-game Lanctot et al. (2017). For simplicity, in this survey, we use the term “restricted game” throughout, regardless of whether payoffs are estimated or not., which acts as an approximation of the underlying full game. Restricted games are induced from simulations run over combinations of a particular set of strategies. This set of strategies is typically much smaller than the full game, and thus restricted games are feasible to analyze with traditional equilibrium computation methods. PSRO iteratively expands the restricted game by introducing new strategies generated via learning, based on the analysis of the current restricted game.
@@
As a general solver for large-scale games, PSRO has been successfully applied to a wide range of game types and diverse application domains, from mechanism design for sequential auctions Zhang et al. (2023) to robust RL Liang et al. (2023). Numerous PSRO variants have been developed, each tailored to leverage the specific characteristics of different underlying games. As a notable example of its success, algorithms inspired by PSRO have reached state-of-the-art performance in large-scale games such as Barrage Stratego McAleer et al. (2020), and in StarCraft Vinyals et al. (2019), where they have convincingly outperformed human experts and prior AI systems.
@@
While PSRO and its variants have been covered to some extent in existing multiagent learning surveys (e.g., Yang and Wang (2020); Long et al. (2023); Albrecht et al. (2024)), a dedicated survey on PSRO like this one has been lacking. In this survey, we reflect on the historical development of PSRO, which arose from different research communities, and we place PSRO within the space of game solving approaches. We then present the latest developments on PSRO, highlighting both current and future research directions.
@@ 2 The PSRO Framework
The PSRO framework incorporates a synthesis of ideas originating from distinct research communities. Within the planning community, McMahan et al. McMahan et al. (2003) laid the groundwork by formulating robust planning in Markov Decision Processes as a zero-sum game characterized by a vast strategy space. Drawing inspiration from Bender’s decomposition in optimization Benders (1962), they introduced the Double Oracle (DO) algorithm. DO is an iterative algorithm for solving games with a finite number of strategies. DO maintains a restricted version of the full game and iteratively expands the restricted game by adding best responses to the current equilibrium. When DO terminates, no player can deviate unilaterally to gain extra payoff and therefore the equilibrium in the current restricted game is an NE of the full game. In finite games, DO is guaranteed to converge to an NE, though the restricted game could include all strategies of the full game in the worst case. Moreover, under proper tie-breaking assumptions, Zhang and Shandholm Zhang and Sandholm (2024) showed that DO takes exponentially many iterations to converge in partially-observable stochastic games and extensive-form games.
@@
Concurrently, similar methods were being developed from a different perspective in the co-evolution research community. Co-evolutionary methods evolve multiple populations of (different) species in parallel. In this setting, there is no explicit fitness function, but the fitness of a population member depends on how well it interacts with members of other populations. Traditional challenges for co-evolution include the presence of intransitive cycles in traits, and the related problem of forgetting a trait that seems not useful at one point in the evolutionary process but later becomes useful again. The community explored memory mechanisms and game formulations where the co-evolutionary process was set up to discover NE as mixtures of traits Angeline et al. (1993); Popovici et al. (2012). To overcome the mentioned obstacles of intransitive cycles and forgetting traits, various archival structures were devised to facilitate monotonic convergence towards a range of game-theoretic solution concepts. For example, the “Nash memory” for symmetric games Ficici and Pollack (2003) identifies equilibrium strategies within a discovered restricted game. This was extended to asymmetric games via the Parallel Nash Memory (PNM) Oliehoek et al. (2006), broadening the standard DO framework to accommodate alternative oracle types beyond the best response.
@@
The PSRO framework Lanctot et al. (2017) was introduced in a paper titled “A Unified Game-Theoretic Approach to Multiagent Reinforcement Learning”, bringing many of the mentioned ideas from the planning, multiagent RL, and co-evolution communities together in a unified framework, which also covers classical game-theoretic learning dynamics such as Fictitious Play Brown (1951). Technically, PSRO generalizes DO in three ways. Firstly, PSRO introduces the concept of meta-strategy solver (MSS), which extracts a profile from the current restricted game as the next best-response target [^2] For simplicity, we often use the same term for the solution concept and the MSS that computes it. For example, we may say “Nash equilibrium” to mean the MSS that computes an NE.. This enables the best-response target to extend beyond NE, transforming PSRO into a versatile framework capable of generalizing various classical and modern game-theoretic algorithms. Secondly, PSRO generalizes DO by allowing any form of (approximate) response oracles, including search, planning, and evolutionary algorithms etc (as used in PNM and other works Oliehoek et al. (2006); Li and Wellman (2021)). This generalization enables best-response computation in environments with a large number of states and actions. Thirdly, compared to DO, the payoffs of profiles in PSRO are estimated through simulation. See Figure 1 for a depiction of the whole framework, contrasted with DO.
We note that the PSRO framework can be thought of as an instance of Empirical Game-Theoretic Analysis (EGTA) Wellman (2006), which includes a broad set of methods that build and analyze restricted games based on simulation. A comprehensive introduction to EGTA can be found in the survey by Wellman et al. Wellman et al. (2024). As a concrete example of the connection between EGTA and PSRO, in an early EGTA work, Schvartzman and Wellman Schvartzman and Wellman (2009) deployed tabular RL as a best-response oracle (at a time when deep RL did not exist yet) and NE of the restricted game as a best-response target for strategy generation (i.e. as the MSS).
@@ 2.1 The Framework
A normal-form (aka strategic-form) representation of the full game $\mathcal{G}=(N,(S_{i}),(u_{i}))$ is a tuple, where $N$ is a finite set of players, each with a non-empty set of strategies $S_{i}$ and a utility function $u_{i}:\Pi_{j\in N}S_{j}\rightarrow\mathbb{R}$. A restricted game $\hat{\mathcal{G}}_{S\downarrow X}=(N,(X_{i}),(\hat{u}_{i}))$ is a projection of the full game $\mathcal{G}$, with players choosing from restricted strategy sets $X_{i}\subseteq S_{i}$, allowing for utilities to be estimated via simulation.
@@ 2.2 Strategy Exploration in PSRO
In essence, game-theoretic analysis in PSRO is performed by reasoning about restricted games. A restricted game is expected to contain an effective subset of strategies for representation tractability yet still represent the full game well strategically Balduzzi et al. (2018). This challenge of restricted game construction with minimum computational cost (i.e., with the fewest strategies required) is described as the strategy exploration problem Jordan et al. (2010), which is the main research focus for developing PSRO methods. In PSRO, strategy exploration can be controlled by setting MSSs and ROs, which have a coupled impact; we refer to the joint choice as an MSS-RO combination.
@@
2.3 RL View of PSRO: Population-Based Training
In PSRO, a restricted game maintains a population of strategies (also known as policies, i.e., RL agents) and expands this population by introducing new strategies that respond to a specific distribution (mixture) of strategies within the existing population. Unlike self-play, which is a standard alternative approach in which a strategy is trained directly against itself, training a new strategy against a diverse set of strategies in the population, as in PSRO, can potentially enhance the robustness of the resulting strategy. Consequently, PSRO can be classified as a variant of population-based training. The textbook by Albrecht et al. Albrecht et al. (2024) provides an excellent explanation of PSRO from this viewpoint.
@@ 2.4 Organization of the Survey
In this survey, we discuss PSRO from the perspective of game-theoretic analysis, that is, how to conduct effective strategy exploration given a specific goal. Although the performance of strategy exploration depends on the interplay between the chosen MSSs and ROs, existing literature predominantly focuses on setting either MSSs or ROs independently. Therefore, we organize our discussion on research directions and corresponding PSRO variants by first discussing setting MSSs and ROs independently (Sections 3 and 4 respectively). We then, in Section 5, discuss works that have investigated the joint choice of MSS-RO combination, before moving on to discuss in Section 6 how best to evaluate the effectiveness of such choices for strategy exploration. We then discuss research on improving the efficiency of PSRO (Section 7), explore applications of PSRO (Section 8), implementations of PSRO (Section 9), and conclude with open questions and future research directions (Section 10).
@@ 3 Strategy Exploration via MSS
In prior works, setting meta-strategy solvers was a primary way to control strategy exploration. In this section, we discuss these works, their motivations, and their efficacy for strategy exploration.
3.1 PSRO with Normal-Form Restricted Games
In the standard PSRO framework, a restricted game is represented in normal form. While the simulations typically unfold through sequential observations and decisions over time, the restricted game abstracts away this temporal structure.
@@ 3.1 PSRO with Normal-Form Restricted Games
In PSRO, the most common MSS target is NE, which can be computed by various game-theoretic methods based on the normal-form restricted game. PSRO with NE is essentially DO with deep RL for computing (approximate) best responses. Therefore, PSRO with NE inherits the convergence property of DO, given mild assumptions about the quality of the best responses: In finite games, as long as beneficial deviations can always be found with non-zero probability, PSRO with NE as MSS target will converge to an NE given enough iterations.
@@
Despite its theoretical convergence guarantee, achieving exact convergence in large games is often unattainable due to constraints such as limited computational resources. Consequently, many prior works studying strategy exploration in large games revolve around the development of new algorithms that exhibit strong empirical performance (e.g., rapid convergence in terms of regret within a small number of PSRO iterations). These works found that a key problem that can prevent good empirical performance is overfitting. For strategy exploration, overfitting can arise in two distinct two ways. First, strategy exploration may overfit to the NE of the restricted game. Due to the limited information in the restricted game, the NE may not be an effective best-response target from a global view (i.e., using it, we may fail to generate non-trivial strategies for full-game play) while other best-response targets in the restricted game could be more effective. The second form of overfitting relates to only capturing a specific equilibrium in general games (e.g., games with more than two players or general-sum games) without sufficiently exploring the whole strategy space. Note that the overfitting can be a problem for any solution concept; we discussed it in the context of NE specifically due to NE’s prominence as the MSS target in the literature.
@@
To address the overfitting problem, Lanctot et al. Lanctot et al. (2017) proposed an MSS, called Projected Replicator Dynamics (PRD), an adaptation of traditional replicator dynamics Taylor and Jonker (1978). PRD ensures a probability lower bound for selecting each strategy in the restricted game, allowing the new best response to train against not only strategies in the equilibrium support but also those outside the support. PRD can be viewed as a form of regularization to prevent overfitting the response to a single exact NE of the restricted game. Due to the diverse training targets, PRD also improves the stability of the new strategy.
Building on the concept of regularization, subsequent research has focused on designing MSSs that prevent overfitting by effectively regularizing the best-response target. For instance, Wang et al. Wang et al. (2019) proposed an MSS that combines NE with a uniform distribution as the best-response target, enabling the best response to an NE strategy mixed with exploration elements. Wright et al. Wright et al. (2019) developed a history-aware approach with a best-response target mixed with previous targets. Online double oracle Dinh et al. (2022) integrated PSRO with online learning and used an online profile as the best-response target, which can be viewed as a form of regularization. Wang and Wellman Wang and Wellman (2023b) and Li et al. Li et al. (2023c) employed quantal response equilibrium McKelvey and Palfrey (1995); Gemp et al. (2022) as an MSS, regularizing with bounded rationality. Wang and Wellman Wang and Wellman (2023b) adopted an explicit view of regularization and introduced Regularized Replicator Dynamics (RRD), an MSS variant that truncates the NE search process in intermediate restricted games based on a regret criterion. Specifically, RRD computes the best-response target by running RD in the restricted game, stopping once the regret of the current profile with respect to the restricted game meets a specified regret threshold. The regret criterion enables RRD to support direct control of the degree of regularization and can be adjusted to fit a specific game.